Suppose the price of gasoline decreases from $4.10 to $3.00, and in response quantity demanded increases from 10700 to 11600. Using the mid-point formula, what is the price elasticity of demand? (Note: your answer should be correct to two decimal places, and remember to take the absolute value of your result.)
Answer #1

Answer

Elasticity of demand=(change in quantity/average quantity)/(change in price/average price)

Change in quantity=3-4.1=-1.1

Average quantity=(3+4.1)/2=3.55

Change in price=11600-10700=900

average price=(11600+10700)/2=11150

Elasticity of demand=(-1.1/3.55)/(900/11150)

=-3.83881064

=-3.84

=3.84 (absolute value)

The elasticity of demand is elastic as the elasticity is above 1.